OpenStudy (ledah):

Are 20, 25 and 30 an example of a Pythagorean triplet?

OpenStudy (shayaan_mustafa):

Hi Ledah :) How are you?

OpenStudy (shayaan_mustafa):

No. It is not the triple. Do you know why?

OpenStudy (ledah):

Not sure.

OpenStudy (ledah):

Was never taught about triplets.

OpenStudy (shayaan_mustafa):

Its OK. I will help you. Don't worry.

OpenStudy (shayaan_mustafa):

Pythagorean Theorem is \[\large H^2=P^2+B^2\] H=hypotenuse P=perpendicular B=base

OpenStudy (shayaan_mustafa):

As you can see all three have square. So no, find in your question that all three are squared of some number or not?

OpenStudy (ledah):

they aren't

OpenStudy (ledah):

It only increases by 5 each time

OpenStudy (shayaan_mustafa):

20= it is not the square of any number, 25= it is the square of 5 30= it is not the square of any number.

OpenStudy (shayaan_mustafa):

Yes, it only increase by 5 each time. So for Pythagorean theorem all three must be squared. Got it?

OpenStudy (ledah):

Yeah I've got it. Thanks

OpenStudy (shayaan_mustafa):

You are ever welcome. Good luck :D